A battery is connected in series with a resistor R. The battery transfers 2000 C of charge completely round the circuit. During this process, 2500 J of energy is dissipated R and 1500 J is expended in the battery. Calculate the emf of the battery. 


1
Expert's answer
2021-12-28T10:56:00-0500

For a battery

E.m.f"=" Total work done in moving a unit charge in a circuit



Total work done"=2500+1500=4000J"



E.m.f "=\frac{4000}{2000}=2V"